OpenStudy (anonymous):

find y'' radical x+ radical y=1

OpenStudy (anonymous):

\[\sqrt{x}+\sqrt{y}=1\]

OpenStudy (anonymous):

\[x^{\frac{ 1 }{ 2 }}+y^{\frac{ 1 }{ 2 }}=1\]

OpenStudy (anonymous):

\[y=(1-x^{\frac{ 1 }{ 2 }})^{2}\]\[y^{'}=2(1-x^{1/2})(-1/2x^{-1/2})\]

OpenStudy (anonymous):

y'=\[-\frac{ \sqrt{x} }{ \sqrt{y}}\]

OpenStudy (anonymous):

opps the other way around

OpenStudy (anonymous):

negative radical y over radical x